You toss a rock up vertically at an initial speed of 40 feet per second and release it at an initial height of 6 feet. The rock will remain in the air for how many
seconds.
It will reach a maximum height of how many
feet after how many
seconds.
Note: Ignore air resistance.
1
Expert's answer
2020-11-29T19:12:38-0500

Let`s start with recognising given valuables:

V0=40 feet/second,

h0=6 feet

And we need to find:

ttotal=?, hmax=?, t1=?

tis time which is spent to reach maximum height.

Firstly, let`s start to compute with the maximum height:

hmax=h0+V02/(2*g) --> g=9.81 m/s2 -->g=32.18 feet/second

we have everything to calculate this and the result is:

hmax=6+402/(2*32.18)=30.86 feet.

Secondly we calculate time to reach maximum height:

t1="\sqrt{\smash[b]{(2*(hmax-h0)/g)}}" =1.24 second

Lastly, we compute the value of total time of flying:

ttotal=t1+"\sqrt{\smash[b]{(2*hmax/g)}}" =2.62 second.

We found everything which are asked.